OpenStudy (anonymous):

lim x^3 * e^(-x/2) x→∞ I have the answer to the problem which is 0 but, I don't understand the derivative of e^(-x/2). can someone explain how that derivative works? I just need to know the derivative, not the answer to the whole problem

OpenStudy (irishboy123):

\[ \left( e^{f(x)}\right)' = f'(x) e^{f(x)}\]

OpenStudy (irishboy123):

make any sense?

OpenStudy (irishboy123):

or see it as \(y = e^u, \ u = f(x)\) then chain rule \(\large \frac{dy}{dx} = \frac {dy}{du}\frac{du}{dx}\)

OpenStudy (anonymous):

ok I see it. thanks again